# ACCESS: Assurance Case Centric Engineering of Safety-critical Systems

### v1

@@ 1 Introduction
Safety-critical systems require justifications that they are acceptably safe to operate in their defined operational contexts. Assurance cases provide an explicit means for arguing, justifying and assessing the confidence in the safety of safety-critical systems. The submission of an assurance case is increasingly being required during system certification in many safety-critical industries, such as aviation eurocontrol , nuclear power iaea , transportation iso26262 ; ukrail , healthcare habli2018safety and defence ukmod . Prior to certification, an assurance case must be rigorously, and often independently, evaluated [^1] In this work, we use evaluation to refer to both validation and verification activities involved in the development and the assessment processes of safety critical systems and their assurance cases. to ensure that the arguments and evidence for safety is coherent and convincing.
@@
Assurance cases are not self-contained documents. They usually depend on a variety of engineering artifacts that provide contextual and evidential information, including requirement documents, architecture designs, behaviour models, safety analyses, etc. These artifacts may originate from diverse languages and tools, and can be used for prototyping, analysis, formal verification, and the derivation of real-world artifacts. Therefore, assurance case evaluation involves the evaluation of the engineering artifacts an assurance case depends on, which is often an informal, manual, and error-prone process. In an idealised development process, an assurance case is the central point of reference for all system stakeholders, to allow effective communication over diverse engineering artifacts. In addition, changes in engineering artifacts require the assurance case to be re-evaluated denney2015dynamic , which can significantly impact development efficiency. This challenge becomes more obvious as systems become more complex. Hence, there is a need to automate some (if not all) of the system assurance activities to efficiently manage assurance cases and their referenced engineering artifacts.
@@
Over the past few years, system assurance practitioners have begun adopting Model Based Systems Engineering (MBSE). MBSE promises the interoperability and management of diverse artifacts/models in an automated manner, which provide the basis for automated, coherent and self-contained assurance cases. However, existing assurance case notations, such as the Goal Structuring Notation (GSN) kelly2004goal and Claim-Argument-Evidence (CAE) bishop2000methodology , do not have a sufficient model-based foundation to systematically support this kind of integration wei2019model . Consequently, existing model-based assurance case approaches cannot provide the collective and automated evaluation of an assurance case, together with the engineering artifacts that it may depend on. The inspection, evaluation, and change management of engineering artifacts still remain manual.
@@
In recent years, new applications for Robotics and Autonomous Systems (RAS) have emerged, which are often safety critical. RASs are increasingly open (they inter-connect at runtime) and adaptive (they adapt to changing contexts at runtime), that render the current generation of safety assurance approaches insufficient trapp2013safety ; denney2015dynamic . Specifically, assurance cases for RAS need to be living documents that can evolve during the operational life of the system with minimal human intervention. As such, it is imperative to shift some system safety assurance activities from development time to runtime trapp2013safety ; wei2018 . This is a significant challenge, though, as it requires automation of verification and validation activities that are both crucial parts of the evaluation process. On the one hand, we must demonstrate that each engineering artifact meets its requirements through verification, and on the other we must ensure that real-world system artifacts exhibit the behaviour predicted by a model through techniques like runtime safety monitoring Machin2018SMOF .
@@
To address the identified challenges, we introduce our model-based, assurance oriented methodology – Assurance Case Centric Engineering of Safety-critical Systems (ACCESS). ACCESS is underpinned by a combination of (1) design-time automated assurance-case-and-engineering-artifacts management and evaluation, and (2) runtime assurance case evaluation based on runtime data. We present a tooling prototype, Assurance Case Management Environment (ACME), which supports the creation and the management of assurance cases based on the Structured Assurance Case Metamodel (SACM) sacm , an international standard.
@@
To demonstrate our approach, we provide a case study on the assurance case for an Autonomous Underwater Vehicle (AUV), including its safety requirements, arguments, and a formal model of the safety controller in the RoboChart language miyazawa2016robochart with formal verification evidence.
@@
We discuss how we can apply our approach to develop critical systems around an evolving assurance case. In addition, we also demonstrate tool support for ACCESS at development time, so that we can: 1) perform a collective evaluation of an assurance case, via model-based traceability, with respect to the engineering artifacts it refers to; 2) automatically invoke evidence from formal methods by using Isabelle/HOL Isabelle as a verification service within an assurance case; 3) automatically generate and machine-check formalisation of an assurance case to verify its logical integrity; and 4) enable automated change impact analysis from engineering artifacts to assurance cases. We also discuss how we can turn a development time assurance case to a dynamic runtime assurance case and discuss how runtime assurance case evaluation (based on runtime data) can be achieved using our approach.
@@
The rest of the paper is organised as follows. In Section 2 we provide some background information on assurance cases, GSN, model-based assurance case and formal methods. In Section 3 we describe the generic ACCESS methodology. In Section LABEL:motivating_example we provide an illustrative example using the ACCESS methodology. In Section 4 we discuss our tool support to back the ACCESS methodology In Section 5 we evaluate the ACCESS methodology in depth with an case study for an Autonomous Underwater Vehicle (AUV). In Section 7 we discuss related work and in Section 8 we conclude this paper with a discussion and point out future research directions.
@@ 2.1 Assurance Cases
The practice of safety certification is increasingly goal-oriented rather than highly prescriptive denney2015dynamic . Goal-oriented certification places greater emphasis on explicitly stating safety claims, and supplying an argument along with supporting evidence to satisfy certification goals that regulators define mcdermid2001software . Examples may include “All identified hazards have been mitigated” or “Necessary assumptions of the physical environment have been defined”. Such arguments and evidence are generally organised in the form of an assurance case. As defined in kelly1999arguing , an assurance case is a document that should communicate a clear, comprehensive and defensible argument that a system is acceptably safe to operate in a particular context.
@@
Conventionally, an assurance case is not a self-contained document. Definitions of assurance case jsp430 ; 0055 indicate that an assurance case is a document (either as a logical concept or as a physical artifact) that can refer to, and pull together information regarding system safety (such as system requirements, system architectural design, safety analyses, etc.), to form a safety argument. The development of an assurance case involves communications among various stakeholders, one typical scenario is illustrated in Figure 1. When a system concept is formed, system engineers define a set of requirements. Based on these requirements, safety engineers may perform safety analyses, from which hazards (and their associated risks) are identified. Identified hazards and risks are then used to derive safety goals, from which safety engineers can elicit safety requirements. Safety requirements are then considered in the system design, which impose constraints (e.g. acceptable failure rates), and mitigation measures (e.g. redundancy and monitoring). As an assurance case is developed, it may refer to all engineering artifacts above within its argument, for contextual and evidential information. Therefore, the evaluation of an assurance case typically involves the validation and verification of engineering artifacts it refers to.
Assurance cases are subject to evolution, where engineering artifacts and arguments need to be adapted, potentially because of new requirements or upgrades. Changes in the engineering artifacts may invalidate the assurance case, and so its integrity must be checked through evaluation. To evaluate an assurance case, practitioners typically need to trace, navigate to, review, validate and verify the engineering artifacts it depends on hawkins2015weaving . This then informs the decision as to whether the system is acceptably safe to deploy and operate in its intended operational context. On the other hand, when an engineering artifact is changed during the development process, its impact in the assurance case needs to be identified, sometimes resulting the assurance case to be re-evaluated. This is observed in nair2015evidence , the authors of which report that evidence completeness and change impact for assurance cases are managed mostly manually using (sometimes even no) traceability information. Importantly, their study raises the question of how evolution and changes are identified, assessed and managed at the level of assurance case.
@@ 2.2 Goal Structuring Notation
In the current state of practice, assurance cases are typically communicated using graphical notations, among which the most widely used notation is the Goal Structuring Notation (GSN) kelly2004goal . GSN is a well established graphical argumentation notation that is widely adopted within safety-critical industries for the presentation of safety arguments within safety cases. The core elements of GSN are shown in Figure 2.
@@
Figure 5: The Assurance Case component of SACM sacm .
@@ 2.3 Model-Based Assurance Cases
Model Based Systems Engineering (MBSE) brambilla2017model is a contemporary systems engineering approach. In MBSE, models are first class artefacts, therefore driving the development. MBSE has been proven to improve consistency and productivity significantly due to the automation provided by model management operations jaaksi2002developing ; karna2009evaluating .
Over the past few years, model-based assurance case approaches emerged due to the benefits introduced by MBSE. Studies have shown how automated MBSE operations can be performed on model-based assurance cases (created using GSN) to check the well-formedness of assurance cases denney2017tool , generate and assemble structured argumentation within assurance cases hawkins2015weaving , and automatically generate texts for assurance case reports denney2017tool . However, existing model-based assurance case approaches (GSN and CAE - Cliams-Arguments-Evidence bishop2000methodology ) do not provide sufficient support for traceability to engineering artifacts. This is partly caused by the fact that GSN and CAE permit only structured arguments and not external artifact traceability. This is a historical problem, as prior to model-based assurance case approaches, GSN and CAE are used to create physical documents, which naturally contain references to other (physical) engineering artifacts by their names.
@@ 2.4 Structured Assurance Case Metamodel
To address this limitation, the Object Management Group (OMG) specified and issued the Structured Assurance Case Metamodel (SACM) sacm . SACM is developed by the specifiers of existing system assurance approaches (e.g. GSN and CAE), based on the collective knowledge and experiences of safety and/or security practitioners over the period of last two decades. Therefore, features that are not previously supported by GSN and CAE have been evaluated and included in SACM.
@@
Figure 6: The Base component of SACM sacm .
@@
Figure 7: The Artifact component of SACM sacm .
@@ 2.5 Formal Methods and RoboChart
Assurance cases can benefit from formal methods Gleirscher2018-NewOpportunitiesIntegrated . Informal safety arguments and evidence can be difficult to automatically evaluate, and may be subject to some argumentation fallacies greenwell2006taxonomy . Thus, formalisation of requirements, to allow the use of formal methods, can significantly improve both automation and the confidence foster2020formal . At the same time, assurance cases allow us to put formal methods results in context. For systems assurance it is never enough to simply prove properties of a formal model. For the results to be meaningful, the model must be linked to its corresponding real-world artifact Lee2018 , such as software and hardware by some form of validation argument. Consequently, a comprehensive demonstration of safety requires both assurance cases and formal methods.
An important development here is Isabelle Isabelle , a verification framework for integrated formal methods Wenzel2007FMIsabelle ; Wenzel2019-Isar ; Foster2019-iFM . Development centres around documents called Isabelle theories, which encode graphs of hyperlinked mathematical artifacts, such as definitions, theorems, and proofs. Formal method integration is supported by (1) a flexible front-end, which supports a variety of languages Tuong2019-CIsabelle and their translation into formal semantics; (2) an extensible plugin-oriented architecture where external tools, such as SMT solvers Blanchette2011 , can improve automation; and (3) incremental theory processing Wenzel2007FMIsabelle . Moreover, Isabelle can be installed as a server component which other tools can make use of as a verification tool service. At the foundational level, Isabelle can be used for mechanisation of a variety of formal semantics Nipkow2014-ConcreteSemantics , to support verification tools. A front-end for a programming language, such as C Tuong2019-CIsabelle , can be developed to support concrete program verification Alkassar2008 . Moreover, SACM was recently implemented in Isabelle to create Isabelle/SACM Foster2019-iFM , which allows verification of assurance cases, integration with formal evidence, and generation of certification documents Brucker2019-DOFCert .
Formal methods can be difficult for non-experts to apply, and so there is a desire to use model-based graphical frontends. For example, the RoboChart language Miyazawa2019-RoboChart is a graphical language for the architectural and behavioural description of a robotic controller. It includes a formalised subset of the UML state machine notation with a complete formal semantics in the CSP process algebra Brookes1984 . The formal semantics allows RoboChart models to be subjected to formal analysis using model checking miyazawa2016robochart ; Miyazawa2019-RoboChart and theorem proving Foster2018-FACS ; foster2020formal . RoboChart is therefore both accessible to practitioners, and at the same time uses formal methods to allow development at higher assurance levels.
@@ 3.4 Step 4: Validate and Verify Engineering Artifacts
With the assurance case model in place, the next step is to investigate how it could be verified and validated, and preferably in an automated manner. In our process, we introduce an additional task, S5: Behavioural Modelling, which is typical for software development process. This task can include the creation of state machines or sequence diagrams for each of the subsystem, conforming to the architectural model. We assume that the behavioural modelling notation will have a formal semantics, as is the case for example with the RoboChart language Miyazawa2019-RoboChart . Next, we propose the task of A5: Model Analysis and Verification. The purpose here is to ensure that each of the requirements is represented in or satisfied by the behavioural model, and reflected in the assurance case. For example, we may wish to state that a particular technical requirement is implemented by a transition, or that a state machine satisfies a global safety requirement. We may also wish to verify the integrity of the model, for example by checking for the possibility of deadlock or livelock using a model checker. This task may expose flaws in the behavioural model, such as inadequately implemented requirements, and so there is iteration between S5 and A5, which should in all circumstances be carried out by separate teams.
@@ 4 Tool Support
The ACCESS methodology is backed by our tool support - Assurance Case Management Environment (ACME). ACME is an integrated model based assurance case modelling and management framework, that supports the creation and management of assurance case models that conform to the Structured Assurance Case Metamodel (SACM) sacm . Since SACM is relatively new and its graphical notations are being standardised, ACME also supports the Goal Structuring Notation (GSN), whose current model-based implementation extends the abstract syntax of SACM, explained in detail in wei2019model . In this way, ACME supports the creation of model-based GSN diagrams, and at the same time provides access to all other features of SACM.
The architecture of ACME is illustrated in Figure 9. We implement SACM and GSN with the Eclipse Modelling Framework (EMF) steinberg2008emf , and use Graphical Modelling Framework (GMF) gmf to create graphical editors for SACM packages and GSN modules.s To enable automated model management and the checking of formal notations, we also integrate:
  • The Eclipse Epsilon platform kolovos2008epsilon , which is an integrated model management platform, that provides task specific model management languages (model validation, model transformations, etc.) that operate on models defined in different modelling technologies (EMF, Excel spreadsheet, Simulink simulink , UML, etc.);
  • Isabelle Isabelle , which is a generic proof assistant that allows mathematical formulas to be expressed in a formal language and provides tools for proving these formulas in a logical calculus, with a high degree of automation. Isabelle also provides an extensible document model which can support the encoding of different meta-models and associated parsers.
@@
To enable the traceability from an assurance case to its supporting engineering artifacts, we make use of SACM’s Artifact component (discussed in Section 2.3) in ACME. With the help of the Artifact component of SACM, in an Artifact element, we are able to record inside it: (1) the type of an engineering artifact (e.g. EMF model, XML document, Excel spreadsheet, Isabelle theory file, etc.) to trace to; (2) the location of the engineering artifact; and (3) the meta-data of the engineering artifact (e.g. metamodel, XML metadata, etc.). To obtain a more fine-grained traceability, we make use of the ImplementationConstraint (discussed in Section 2.3) element, and record model querying/validation programs written in a model querying language (in this work we support the Epsilon Object Language (EOL) kolovos2006epsilon , but any other languages can be supported) in Artifact model elements. In this way, when the program is executed, we are able to obtain specific model element(s) (or information) from the engineering artifact that an Artifact element refers to, which can be used in assurance cases to provide contextual and evidential information. In summary, traceability from an assurance case to external engineering artifacts is achieved by referring to Artifact elements (organised in an ArtifactPackage) that contain traces to engineering models, from (in GSN terms) either an Context or an Solution element organised in a GSN Module.
@@
With the traceability from an assurance case to engineering artifacts, we are also able to perform automated assurance case evaluations (i.e. validation and verification on referenced external engineering artifacts). For validation, we refer to external engineering artifacts using model elements defined in SACM’s Artifact component, and embed programs such as validation rules (which return true or false). ACME executes the validation rules and reflects the validation results to ACME editors so that the users can find out which part of an assurance case failed in the evaluation. For formal verification, we refer to Isabelle theory documentss. A theory document is a hierarchical structure consisting of formal artifacts, such as data types, functions, theorems, and proofs. Upon execution, ACME sends the theory files to the established Isabelle server (discussed in Wenzel2019-Isar ), which can be communicated with using an RESTful API. When a theory file is sent to the Isabelle server, the server processes the file and returns JSON messages conveying the status of all artifacts contained within the theory file. If the processing of any artifacts fail, the JSON messages contain all the problems that Isabelle found. For example, a candidate proof could fail to prove a theorem, and this would raise an error. In ACME, we trace to an Isabelle theory file with an Artifact and perform formal notation checking in an automated manner. If an Artifact cannot be verified, ACME reflects this information in the model editor. Moreover, ACME also supports the translation of an assurance case to an Isabelle/SACM theory Foster2019-iFM , which can be used for verifying its logical integrity. If there are any errors in other assurance case nodes, these are likewise reflected.
@@
In ACCESS Step 5, we perform the evaluation of the entire assurance case, which includes the verification and validation on the system level. Using ACME, we perform two types of evaluation. The first type is to invoke evaluation on all referenced engineering artifacts from the assurance case. In ACME we provide an evaluate function which can be called on an assurance case, ACME then automatically looks into all ArgumentPackages and looks for argument elements that refer to artifact elements in ArtifactPackages, then ACME automatically calls Isabelle or Epsilon and determines if all evidence (that support the assurance case) are valid. The second type is to formalise the assurance case and check its logical integrity in an automated manner. We do this by transforming the assurance case to Isabelle/SACM notations with ACME’s built-in model-to-text transformation (written in the Epsilon Generation Language rose2008epsilon – EGL) and generate the Isabelle/SACM notation representation of the assurance case in the form of a theory file. The Isabelle/SACM notation is then sent to the Isabelle server for machine-checking. The evaluation result will be parsed by ACME, which then locates the model elements in the assurance case that fail the evaluation, and displays their corresponding error messages.

### v2

@@ 1 Introduction
Safety-critical systems require justifications that they are acceptably safe to operate in their defined operational contexts. Assurance cases provide an explicit means for arguing, justifying and assessing the confidence in the safety of safety-critical systems. The submission of an assurance case is increasingly being required during system certification in many safety-critical industries, such as aviation [17], nuclear power [27], transportation [28, 60], healthcare [25] and defence [59]. Prior to certification, an assurance case must be rigorously, and often independently, evaluated [^1] In this work, we use evaluation to refer to both validation and verification activities involved in the development and the assessment processes of safety critical systems and their assurance cases. to ensure that the arguments and evidence for safety is coherent and convincing.
@@
Assurance cases are not self-contained documents. They usually depend on a variety of engineering artifacts that provide contextual and evidential information, including requirement documents, architecture designs, behaviour models, safety analyses, etc. These artifacts may originate from diverse languages and tools, and can be used for prototyping, analysis, formal verification, and the derivation of real-world artifacts. Therefore, assurance case evaluation involves the evaluation of the engineering artifacts an assurance case depends on, which is often an informal, manual, and error-prone process. In an idealised development process, an assurance case is the central point of reference for all system stakeholders, to allow effective communication over diverse engineering artifacts. In addition, changes in engineering artifacts require the assurance case to be re-evaluated [14], which can significantly impact development efficiency. This challenge becomes more obvious as systems become more complex. Hence, there is a need to automate some (if not all) of the system assurance activities to efficiently manage assurance cases and their referenced engineering artifacts.
@@
Over the past few years, system assurance practitioners have begun adopting Model Based Systems Engineering (MBSE). MBSE promises the interoperability and management of diverse artifacts/models in an automated manner, which provide the basis for automated, coherent and self-contained assurance cases. However, existing assurance case notations, such as the Goal Structuring Notation (GSN) [31] and Claim-Argument-Evidence (CAE) [5], do not have a sufficient model-based foundation to systematically support this kind of integration [65]. Consequently, existing model-based assurance case approaches cannot provide the collective and automated evaluation of an assurance case, together with the engineering artifacts that it may depend on. The inspection, evaluation, and change management of engineering artifacts still remain manual.
@@
In recent years, new applications for Robotics and Autonomous Systems (RAS) have emerged, which are often safety critical. RASs are increasingly open (they inter-connect at runtime) and adaptive (they adapt to changing contexts at runtime), that render the current generation of safety assurance approaches insufficient [55, 14]. Specifically, assurance cases for RAS need to be living documents that can evolve during the operational life of the system with minimal human intervention. As such, it is imperative to shift some system safety assurance activities from development time to runtime [55, 67]. This is a significant challenge, though, as it requires automation of verification and validation activities that are both crucial parts of the evaluation process. On the one hand, we must demonstrate that each engineering artifact meets its requirements through verification, and on the other we must ensure that real-world system artifacts exhibit the behaviour predicted by a model through techniques like runtime safety monitoring [38].
@@
To address the identified challenges, we introduce our model-based, assurance oriented methodology – Assurance Case Centric Engineering of Safety-critical Systems (ACCESS). ACCESS is underpinned by a combination of (1) design-time automated assurance-case-and-engineering-artifacts management and evaluation, and (2) runtime assurance case evaluation based on runtime data. We present a tooling prototype, Assurance Case Management Environment (ACME), which supports the creation and the management of assurance cases based on the Structured Assurance Case Metamodel (SACM) [48], an international standard.
@@
To demonstrate our approach, we provide a case study on the assurance case for an Autonomous Underwater Vehicle (AUV), including its safety requirements, arguments, and a formal model of the safety controller in the RoboChart language [43] with formal verification evidence.
@@
We discuss how we can apply our approach to develop critical systems around an evolving assurance case. In addition, we also demonstrate tool support for ACCESS at development time, so that we can: 1) perform a collective evaluation of an assurance case, via model-based traceability, with respect to the engineering artifacts it refers to; 2) automatically invoke evidence from formal methods by using Isabelle/HOL [47] as a verification service within an assurance case; 3) automatically generate and machine-check formalisation of an assurance case to verify its logical integrity; and 4) enable automated change impact analysis from engineering artifacts to assurance cases. We also discuss how we can turn a development time assurance case to a dynamic runtime assurance case and discuss how runtime assurance case evaluation (based on runtime data) can be achieved using our approach.
@@
The rest of the paper is organised as follows. In Section 2 we provide some background information on assurance cases, GSN, model-based assurance case and formal methods. In Section 3 we describe the generic ACCESS methodology. In Section 4 we discuss our tool support to back the ACCESS methodology In Section 5 we evaluate the ACCESS methodology in depth with an case study for an Autonomous Underwater Vehicle (AUV). In Section 7 we discuss related work and in Section 8 we conclude this paper with a discussion and point out future research directions.
@@ 2.1 Assurance Cases
The practice of safety certification is increasingly goal-oriented rather than highly prescriptive [14]. Goal-oriented certification places greater emphasis on explicitly stating safety claims, and supplying an argument along with supporting evidence to satisfy certification goals that regulators define [41]. Examples may include “All identified hazards have been mitigated” or “Necessary assumptions of the physical environment have been defined”. Such arguments and evidence are generally organised in the form of an assurance case. As defined in [32], an assurance case is a document that should communicate a clear, comprehensive and defensible argument that a system is acceptably safe to operate in a particular context.
@@
Conventionally, an assurance case is not a self-contained document. Definitions of assurance case [57, 58] indicate that an assurance case is a document (either as a logical concept or as a physical artifact) that can refer to, and pull together information regarding system safety (such as system requirements, system architectural design, safety analyses, etc.), to form a safety argument. The development of an assurance case involves communications among various stakeholders, one typical scenario is illustrated in Figure 1. When a system concept is formed, system engineers define a set of requirements. Based on these requirements, safety engineers may perform safety analyses, from which hazards (and their associated risks) are identified. Identified hazards and risks are then used to derive safety goals, from which safety engineers can elicit safety requirements. Safety requirements are then considered in the system design, which impose constraints (e.g. acceptable failure rates), and mitigation measures (e.g. redundancy and monitoring). As an assurance case is developed, it may refer to all engineering artifacts above within its argument, for contextual and evidential information. Therefore, the evaluation of an assurance case typically involves the validation and verification of engineering artifacts it refers to.
Assurance cases are subject to evolution, where engineering artifacts and arguments need to be adapted, potentially because of new requirements or upgrades. Changes in the engineering artifacts may invalidate the assurance case, and so its integrity must be checked through evaluation. To evaluate an assurance case, practitioners typically need to trace, navigate to, review, validate and verify the engineering artifacts it depends on [26]. This then informs the decision as to whether the system is acceptably safe to deploy and operate in its intended operational context. On the other hand, when an engineering artifact is changed during the development process, its impact in the assurance case needs to be identified, sometimes resulting the assurance case to be re-evaluated. This is observed in [44], the authors of which report that evidence completeness and change impact for assurance cases are managed mostly manually using (sometimes even no) traceability information. Importantly, their study raises the question of how evolution and changes are identified, assessed and managed at the level of assurance case.
@@ 2.2 Goal Structuring Notation
In the current state of practice, assurance cases are typically communicated using graphical notations, among which the most widely used notation is the Goal Structuring Notation (GSN) [31]. GSN is a well established graphical argumentation notation that is widely adopted within safety-critical industries for the presentation of safety arguments within safety cases. The core elements of GSN are shown in Figure 2.
@@
Figure 5: The Assurance Case component of SACM [48].
@@ 2.3 Model-Based Assurance Cases
Model Based Systems Engineering (MBSE) [7] is a contemporary systems engineering approach. In MBSE, models are first class artefacts, therefore driving the development. MBSE has been proven to improve consistency and productivity significantly due to the automation provided by model management operations [29, 30].
Over the past few years, model-based assurance case approaches emerged due to the benefits introduced by MBSE. Studies have shown how automated MBSE operations can be performed on model-based assurance cases (created using GSN) to check the well-formedness of assurance cases [13], generate and assemble structured argumentation within assurance cases [26], and automatically generate texts for assurance case reports [13]. However, existing model-based assurance case approaches (GSN and CAE - Cliams-Arguments-Evidence [5]) do not provide sufficient support for traceability to engineering artifacts. This is partly caused by the fact that GSN and CAE permit only structured arguments and not external artifact traceability. This is a historical problem, as prior to model-based assurance case approaches, GSN and CAE are used to create physical documents, which naturally contain references to other (physical) engineering artifacts by their names.
@@ 2.4 Structured Assurance Case Metamodel
To address this limitation, the Object Management Group (OMG) specified and issued the Structured Assurance Case Metamodel (SACM) [48]. SACM is developed by the specifiers of existing system assurance approaches (e.g. GSN and CAE), based on the collective knowledge and experiences of safety and/or security practitioners over the period of last two decades. Therefore, features that are not previously supported by GSN and CAE have been evaluated and included in SACM.
@@
Figure 6: The Base component of SACM [48].
@@
Figure 7: The Artifact component of SACM [48].
@@ 2.5 Formal Methods and RoboChart
Assurance cases can benefit from formal methods [23]. Informal safety arguments and evidence can be difficult to automatically evaluate, and may be subject to some argumentation fallacies [24]. Thus, formalisation of requirements, to allow the use of formal methods, can significantly improve both automation and the confidence [20]. At the same time, assurance cases allow us to put formal methods results in context. For systems assurance it is never enough to simply prove properties of a formal model. For the results to be meaningful, the model must be linked to its corresponding real-world artifact [36], such as software and hardware by some form of validation argument. Consequently, a comprehensive demonstration of safety requires both assurance cases and formal methods.
An important development here is Isabelle [47], a verification framework for integrated formal methods [69, 68, 19]. Development centres around documents called Isabelle theories, which encode graphs of hyperlinked mathematical artifacts, such as definitions, theorems, and proofs. Formal method integration is supported by (1) a flexible front-end, which supports a variety of languages [56] and their translation into formal semantics; (2) an extensible plugin-oriented architecture where external tools, such as SMT solvers [6], can improve automation; and (3) incremental theory processing [69]. Moreover, Isabelle can be installed as a server component which other tools can make use of as a verification tool service. At the foundational level, Isabelle can be used for mechanisation of a variety of formal semantics [46], to support verification tools. A front-end for a programming language, such as C [56], can be developed to support concrete program verification [1]. Moreover, SACM was recently implemented in Isabelle to create Isabelle/SACM [19], which allows verification of assurance cases, integration with formal evidence, and generation of certification documents [9].
Formal methods can be difficult for non-experts to apply, and so there is a desire to use model-based graphical frontends. For example, the RoboChart language [43] is a graphical language for the architectural and behavioural description of a robotic controller. It includes a formalised subset of the UML state machine notation with a complete formal semantics in the CSP process algebra [8]. The formal semantics allows RoboChart models to be subjected to formal analysis using model checking [42, 43] and theorem proving [18, 20]. RoboChart is therefore both accessible to practitioners, and at the same time uses formal methods to allow development at higher assurance levels.
@@ 3.4 Step 4: Validate and Verify Engineering Artifacts
With the assurance case model in place, the next step is to investigate how it could be verified and validated, and preferably in an automated manner. In our process, we introduce an additional task, S5: Behavioural Modelling, which is typical for software development process. This task can include the creation of state machines or sequence diagrams for each of the subsystem, conforming to the architectural model. We assume that the behavioural modelling notation will have a formal semantics, as is the case for example with the RoboChart language [43]. Next, we propose the task of A5: Model Analysis and Verification. The purpose here is to ensure that each of the requirements is represented in or satisfied by the behavioural model, and reflected in the assurance case. For example, we may wish to state that a particular technical requirement is implemented by a transition, or that a state machine satisfies a global safety requirement. We may also wish to verify the integrity of the model, for example by checking for the possibility of deadlock or livelock using a model checker. This task may expose flaws in the behavioural model, such as inadequately implemented requirements, and so there is iteration between S5 and A5, which should in all circumstances be carried out by separate teams.
@@ 4 Tool Support
The ACCESS methodology is backed by our tool support - Assurance Case Management Environment (ACME). ACME is an integrated model based assurance case modelling and management framework, that supports the creation and management of assurance case models that conform to the Structured Assurance Case Metamodel (SACM) [48]. Since SACM is relatively new and its graphical notations are being standardised, ACME also supports the Goal Structuring Notation (GSN), whose current model-based implementation extends the abstract syntax of SACM, explained in detail in [65]. In this way, ACME supports the creation of model-based GSN diagrams, and at the same time provides access to all other features of SACM.
The architecture of ACME is illustrated in Figure 9. We implement SACM and GSN with the Eclipse Modelling Framework (EMF) [54], and use Graphical Modelling Framework (GMF) [16] to create graphical editors for SACM packages and GSN modules.s To enable automated model management and the checking of formal notations, we also integrate:
  • The Eclipse Epsilon platform [34], which is an integrated model management platform, that provides task specific model management languages (model validation, model transformations, etc.) that operate on models defined in different modelling technologies (EMF, Excel spreadsheet, Simulink [39], UML, etc.);
  • Isabelle [47], which is a generic proof assistant that allows mathematical formulas to be expressed in a formal language and provides tools for proving these formulas in a logical calculus, with a high degree of automation. Isabelle also provides an extensible document model which can support the encoding of different meta-models and associated parsers.
@@
To enable the traceability from an assurance case to its supporting engineering artifacts, we make use of SACM’s Artifact component (discussed in Section 2.3) in ACME. With the help of the Artifact component of SACM, in an Artifact element, we are able to record inside it: (1) the type of an engineering artifact (e.g. EMF model, XML document, Excel spreadsheet, Isabelle theory file, etc.) to trace to; (2) the location of the engineering artifact; and (3) the meta-data of the engineering artifact (e.g. metamodel, XML metadata, etc.). To obtain a more fine-grained traceability, we make use of the ImplementationConstraint (discussed in Section 2.3) element, and record model querying/validation programs written in a model querying language (in this work we support the Epsilon Object Language (EOL) [33], but any other languages can be supported) in Artifact model elements. In this way, when the program is executed, we are able to obtain specific model element(s) (or information) from the engineering artifact that an Artifact element refers to, which can be used in assurance cases to provide contextual and evidential information. In summary, traceability from an assurance case to external engineering artifacts is achieved by referring to Artifact elements (organised in an ArtifactPackage) that contain traces to engineering models, from (in GSN terms) either an Context or an Solution element organised in a GSN Module.
@@
With the traceability from an assurance case to engineering artifacts, we are also able to perform automated assurance case evaluations (i.e. validation and verification on referenced external engineering artifacts). For validation, we refer to external engineering artifacts using model elements defined in SACM’s Artifact component, and embed programs such as validation rules (which return true or false). ACME executes the validation rules and reflects the validation results to ACME editors so that the users can find out which part of an assurance case failed in the evaluation. For formal verification, we refer to Isabelle theory documentss. A theory document is a hierarchical structure consisting of formal artifacts, such as data types, functions, theorems, and proofs. Upon execution, ACME sends the theory files to the established Isabelle server (discussed in [68]), which can be communicated with using an RESTful API. When a theory file is sent to the Isabelle server, the server processes the file and returns JSON messages conveying the status of all artifacts contained within the theory file. If the processing of any artifacts fail, the JSON messages contain all the problems that Isabelle found. For example, a candidate proof could fail to prove a theorem, and this would raise an error. In ACME, we trace to an Isabelle theory file with an Artifact and perform formal notation checking in an automated manner. If an Artifact cannot be verified, ACME reflects this information in the model editor. Moreover, ACME also supports the translation of an assurance case to an Isabelle/SACM theory [19], which can be used for verifying its logical integrity. If there are any errors in other assurance case nodes, these are likewise reflected.
@@
In ACCESS Step 5, we perform the evaluation of the entire assurance case, which includes the verification and validation on the system level. Using ACME, we perform two types of evaluation. The first type is to invoke evaluation on all referenced engineering artifacts from the assurance case. In ACME we provide an evaluate function which can be called on an assurance case, ACME then automatically looks into all ArgumentPackages and looks for argument elements that refer to artifact elements in ArtifactPackages, then ACME automatically calls Isabelle or Epsilon and determines if all evidence (that support the assurance case) are valid. The second type is to formalise the assurance case and check its logical integrity in an automated manner. We do this by transforming the assurance case to Isabelle/SACM notations with ACME’s built-in model-to-text transformation (written in the Epsilon Generation Language [51] – EGL) and generate the Isabelle/SACM notation representation of the assurance case in the form of a theory file. The Isabelle/SACM notation is then sent to the Isabelle server for machine-checking. The evaluation result will be parsed by ACME, which then locates the model elements in the assurance case that fail the evaluation, and displays their corresponding error messages.